OpenStudy (anonymous):

look at the trinomal : 100k^4+55k^2-75k write the polynimal as the product of the gcf of all its terms and a polynomial

OpenStudy (anonymous):

100 = 2 * 2 * 5 * 5 55 = 5 * 11 75 = 5 * 5 * 3 they all have one 5 in common so you pull out a 5 they all have 1 "k" in common so you pull out a "k" 5k(20k^3 + 11k - 15)